OpenStudy (anonymous):

...an auto part has 95% chance of being made within its tolerance level and a 5% chance of bieng pulled as defective.what is the probability that in a box of 8 parts,no more than 1 is defective. a.9.43 b.0.66 c.0.94 d.0.28

OpenStudy (helder_edwin):

say p=5% is the probability of being defective, then q=1-p=95% is the probability of being not defective. u have n=8 (8 parts in total) then \[ \large P(X=x)=\binom{n}{x}p^xq^{n-x} \] does it make any sense with what u r studying??